OpenStudy (anonymous):

prove the following derivative of hyperbolic function 1.d/dx(sinhx)=1/2(e^x-e^xd/dx(-x)) 2.d/dx(coshx)=1/2(e^x e^-x)-coshx

OpenStudy (anonymous):

cosh(x) = (e^(x) + e^(-x))/2 sinh(x) = (e^(x) - e^(-x))/2 tanh(x) = (e^(x) - e^(-x)) / (e^(x) + e^(-x)) Split the fractions and differentiate as normal d/dx of e^(ux) = ue^(ux)

OpenStudy (anonymous):

So for cosh(x): d/dx[ (e^(x))/2 + e^(-x)/2 ] (1/2)(e^x) + (-1/2)(e^-x) Or, (e^(x) - e^(-x))/2 dx